Two vertices of a triangle at (3,2),(-2,1) and its centroid is (5/3_-1/3). Find third vertex

Grade:10

3 Answers

Amreen
7 years ago
centroid(5/3,-1/3)
let the third vertex be(x,y)
centroid  G=(x1+x2+x3)/3,(y1+y2+y3)/3
acc.to this 5/3=(3+(-2)+x)/3
on solving for x, we get x=4
similarly
(2+1+y)/3=-1/3
so, y=-4
thus vertex is(4,-4)
Nishi
5 years ago
 X1=3, X2=-2,X3=X.  Y1=2,y2=1, y3=y .                         Centroid=5/3 , -1/3
 5/3=3+(-2)+X/3.     ;  -1/3 =2+1+y/3.                            5=3-2+x.                 ; -1=3+y.                                           5=1+X .                  ;.  -1= 3+y.                                          X=4 .                     ;. Y=-4
ankit singh
3 years ago
(x,y)=(\dfrac{x_1+x_2+x_3}{3},\dfrac{y_1+y_2+y_3}{3})
(\dfrac{5}{3},-\dfrac{1}{3})=(\dfrac{3-2+a}{3},\dfrac{2+1+b}{3})
\dfrac{5}{3}=\dfrac{a+1}{3}\text{ and }-\dfrac{1}{3}=\dfrac{3+b}{3}
a=4\text{ and }b=-4
Hence, The coordinate of C:(4,-4)